OpenStudy (anonymous):

Consider an RL circuit with R=10 ohms, L=10 microHenry, and V = 30 volts. Suppose that I = 0 at t=0. Find the energy stored in the inductor as t goes to infinity.

OpenStudy (anonymous):

It has been a while since I played with induction problems, so I may be a little rusty. The instantaneous power supplied to an inductor is given by \[P=iv=L\frac{di}{dt}\] and hence the energy \(E\) stored on an inductor after time \(t\) is given by \[E=\int\limits_{o}^{t}Pdt=\int\limits_{0}^{I}Li di=\frac{1}{2}LI^2\], where \(I^2\) is given by \[I^2=\frac{V^2}{R^2}\] Hence \[E=\frac{LV^2}{2R^2}=\frac{(10\times10^{-6})(30)^2}{2(10)^2}=4.5\times10^{-5}\rm{J}\]or 45 \(\mu\)J. I think this is correct, but as I said, I'm rusty, so you may want to get a second opinion.

OpenStudy (anonymous):

its 1/2 Li^2 i = V/R there fore energy = 0.5 X 10^-2 X 9 = 4.5 X 10^-2